OpenStudy (logan13):

If the diameter of the circle above is 16 mm, what is the area of the circle? 16 mm2 8 mm2 256 mm2 64 mm2

OpenStudy (alexandervonhumboldt2):

radius=1/2*diameter area=radius^2*pi

OpenStudy (logan13):

i think c

OpenStudy (alexandervonhumboldt2):

but i didnt gave answers? 64 pi or 200.96

OpenStudy (alexandervonhumboldt2):

yes C

OpenStudy (alexandervonhumboldt2):

i wasnt gvaing answer B

OpenStudy (logan13):

d

OpenStudy (logan13):

it was d

OpenStudy (unanimoose):

Your options probably say like: \[64\pi mm^2\]

OpenStudy (unanimoose):

Just you can't paste the pi symbol, so most of the times you don't have to actually input pi and you just need radius^2 with the pi symbol next to it.

OpenStudy (unanimoose):

So the radius is half of the diameter. 16/2 = 8 8^2 = 64 64pi = answer like you already saw